### v1

@@ 1 Introduction
Sustainability, intended as having a non-detrimental impact on the environment, society, and the economy, is becoming increasingly essential for humanity’s future. Various efforts are being undertaken at different levels of the societal and economic hierarchy, from country-level to institutional- and business-level entities, and in some cases, down to individuals. These efforts are focused on making these entities sustainable, amid reputational risks NIKOLAOU2015499 , and pressing challenges including global climate change, widespread inequity, governance malpractices, and geopolitical instability folque2021sustainable .
In this work, we focus on sustainability analyses of entities that provide sustainability reports (also referred to as sustainability disclosures); in particular public companies and institutions. These entities are crucial components in the global development of sustainability corporatesusimpt , and we refer to their sustainability as corporate sustainability.
Given these entities’ substantial stake in global sustainability, the analysis of their sustainability, or corporate sustainability analysis, is critical rajesh2020exploring . Yet, this is extremely challenging due to the inherent complexity of corporate sustainability as a concept. To elaborate, the complexities within firms (e.g., geography, size, business activities) and outside the firms (e.g., their global supply chains) as well as the firms’ relationships with non-business stakeholders lead to the evolving frameworks and guidelines put forward to address corporate sustainability issues buallay2020sustainability derqui2020towards . Further complicating the matter is the lack of globally mandated standards for sustainability reporting that guides companies’ efforts in determining the extent and nature of their sustainability disclosures esgdivergence .
For example, a firm may opt to use a convenient corporate sustainability reporting framework that magnifies only certain aspects of a firm’s practices Darnall2022 , while omitting key sustainability dimensions demastus2024organizational . Or, within a given regulatory disclosure framework that a firm shall abide to, it may provide sustainability disclosures that are difficult for external parties to digest, and whose integrity and transparency may be questionable Boiral2015 .
@@
Indeed, today’s corporate sustainability analysis is carried out by human analysts who read sustainability disclosures provided by companies (i.e., the data dimension), and provide corporate sustainability evaluations based on a combination of sustainability frameworks (e.g., Environmental, Social, Governance (ESG) Dmuchowski2023 , Global Reporting Initiative (GRI) Machado2021 , Sustainability Accounting Standards Board (SASB) Eng2022 , Greenhouse Gas Protocol and Carbon Disclosure Project (CDP) Pitrakkos2020 , United Nations’ Sustainable Development Goals (SDG) pradhan2017systematic , United Nation Global Compact (UNGC) orzes2018united ). These evaluations require an extensive amount of human hours Ni2023 , and yet are inevitably influenced by the inherent biases in the frameworks adopted as well as the analysts’ own subjectivity buyukozkan2018sustainability anlaystbias .
@@
We argue that, to address the subjectivity inherent to data and analysts, it is crucial to adopt artificial intelligence (AI), specifically through linguistic understanding methods enhanced by explainable AI (XAI) capabilities, or what is termed as explainable natural language processing (XNLP) xnlp . To qualify, adopting XNLP entails complementing human analysts instead of replacing them. Within corporate sustainability analysis, analysts should remain ‘in the loop’ – while XNLP can expeditiously process numerous sustainability disclosures to unveil consistent and meaningful insights, analysts can further synergise these insights to inform their corporate sustainability assessments.
@@ 2.1 The data dimension
Data is the first building block, or entry point, of corporate sustainability analysis, as depicted in (Fig. 1(A)). While the data required for corporate sustainability assessments can come from various sources, corporate sustainability disclosures are a predominant source nguyen2020empirical . These disclosures consist of reports released by businesses detailing their own sustainability efforts nguyen2020empirical . For example, general sustainability reports typically follow established sustainability frameworks to describe a company’s sustainability efforts and achievements abeysekera2022framework , and are typically accompanied by integrated reports that aggregate the general sustainability information with financial implications fasan2013annual .
@@
Additional data may come from government-mandated sustainability disclosures, where companies are required to release specific aspects of their sustainability performance tan2022assembling , such as their environmental impact Scott2022 , or ESG (Environmental, Social, and Governance) practices SECsusdiclosures . Data originating from NGO sources and media coverage may also be relevant to assess a company’s sustainability efforts WONG2022101045 ngoimpactinfo . Yet these data are influenced by sustainability disclosures materialityanalysismedia2020 fiaschi2020bad , underlining the importance of focusing our discussion on the sustainability disclosures themselves.
@@
Due to the intricacies and multi-faceted dimensions of corporate sustainability, these data sources inherently contain subjectivity. They involve competing notions of corporate sustainability, expressed through a vast and diverse array of sustainability frameworks complexitydiffframeworkruiz Meuer2020 . Moreover, even within a specific framework, deciding what information to convey necessitates a value judgement (e.g., financial vs. environmental materiality), further exacerbating the partiality involved beske2020materiality . We argue that this subjectivity at the data level manifests itself in four data grand challenges that analysts need to face, namely incompleteness, unreliability, ambiguity, and sophistication. These 4 data challenges are reported on the first block of (Fig. 1(A)), referred to as data dimension. We detail them in the following.
@@
Incompleteness. It refers to how data may not provide a comprehensive and complete view of a firm’s sustainability efforts. As certain sustainability disclosures are voluntary, companies can decide on what is material enough to release as sustainability information, thereby raising concerns about openness and the omission of important data beske2020materiality .
@@
Unreliability. It refers to the trustworthiness and accuracy of sustainability data. Indeed, corporate and government-linked sustainability disclosures may involve greenwashing, where organisations misleadingly claim and exaggerate that their activities or products are more environmentally friendly than they really are de2020concepts . Or they use non-transparent carbon offset products, that often do not deliver the offsets promised delacote2024strong . In fact, such developments are particularly frequent for firms that are of larger sizes, to contend with increased dealing with stakeholders greenwasharticle .
@@
Ambiguity. It refers to data being vague and unclear, making negative content less conspicuous RePEc:inm:ororsc:v:30:y:2019:i:6:p:1207-1231 . One example is sharing information without accompanying it with the appropriate context higgins2020managing . This may lead analysts to misunderstand the sustainability efforts of companies. Ambiguity is related, to some extent, to greenwashing zharfpeykan2021representative , although it does not explicitly refer to inaccurate or misleading data as unreliability does.
@@
Sophistication. It refers to sustainability disclosures being tediously lengthy, text-heavy, and requiring specialised knowledge to comprehend, making them an extremely sophisticated writing category smeuninx2020measuring . As a result, firm sustainability reports can require many hours of human effort to understand, let alone derive useful insights goel2020mining .
@@ 2.2 The analyst dimension
Human analysts must synthesise and interpret sustainability data to produce accurate corporate sustainability assessments. To describe this process, relevant data is collected according to the objectives and scope of the corporate sustainability analysis (i.e. evaluating all operations versus one area). Thereafter, the analysis entails leveraging sustainability frameworks or metrics (i.e., carbon footprint) to measure different impact dimensions (environmental, social, economic etc.) tsalis2020new . The costs and inaccuracies of this process are exacerbated by human limitations. Specifically, human analysts cannot deal well with the challenges of subjective data, and are not free from subjective interpretations. These human-centric issues are encapsulated by limited human-hours and potential bias.
Limited human-hours. Analysts have a limited amount of time to produce corporate sustainability assessments. However, the process entails extracting meaningful information from corporate sustainability disclosures goel2020mining , which requires a significant amount of time i.e., human-hours Ni2023 . In particular, the disclosures are extremely time consuming to read and understand due to the sophistication and ambiguity of the data within them. Moreover, given the potential incompleteness of disclosure information, analysts may analyse additional sources to ascertain a firm’s sustainability, further increasing the human-hours required qorri2022practical .
Potential Bias. Analysts’ interpretations of sustainability data are subject to bias. This is not least due to the complex and wide ranging factors to be considered, such as firm characteristics, evolving sustainability guidelines, and economic implications, amongst others, often leading to leading to varying interpretations amongst analysts buyukozkan2018sustainability anlaystbias . To further compound the potential for bias, data ambiguity and unreliability hinders clear interpretation of sustainability information by obfuscating negative content RePEc:inm:ororsc:v:30:y:2019:i:6:p:1207-1231 .
@@ 2.3 NLP to the rescue
Subjectivity pervades through the data and analyst dimensions, making corporate sustainability analyses an onerous task. To alleviate this burden, we argue that AI technologies, through linguistic understanding enabled by natural language processing (NLP), can significantly aid analysts. NLP can automatically process hundreds of sustainability disclosures stammbach2023environmental , summarising key details and extracting insights for analysts. This reduces the human hours required to analyse disclosures and partially addresses potential bias by providing more consistent and replicable insights abram2020methods Chowdhary2020NLP . We concretise NLP’s usefulness in this domain by detailing five NLP tasks for corporate sustainability analysis (T1-T5) that we deem critical, and that have been partially explored in the literature.
@@
(T1) Topic extraction. Topic extraction allows an analyst to obtain, from sustainability disclosures, content most pertinent to corporate sustainability. Consequently, analysts do not need to sieve through all information, and can focus their analysis on material information BINGLER2022102776 . Thematic extraction involves classifying sentences within sustainability frameworks. This includes sorting text into GRI topics polignano-etal-2022-nlp , ESG pillars LEE2023119726 , or ESG-related concepts kang-el-maarouf-2022-finsim4 . Universal extraction filters textual data into generic categories that are separate from sustainability frameworks. For instance, labelling environmental claims stammbach2023environmental , emissions targets ClimateBERT-NetZero , or climate relevant information BINGLER2022102776 . On the other hand, topic discovery determines topics from sustainability disclosures a posteriori. For example, sustainability topics were uncovered from the sustainability disclosures of shipping companies Zhou2021 .
@@
(T2) Sentiment Classification. Sentiment insights can also be derived from sustainability disclosures, providing analysts a valuable perspective on corporate sustainability. Polarity classification determines the emotional polarity of text from sustainability disclosures. For instance, within SEC 10k reports, sustainability related sentences can be labelled as positive, negative or neutral finetuneTransformerESG . Alternatively, strategic classification uncovers a text’s prospective impact. Examples include the risk and opportunity impact analysis of ESG-related texts tseng2023dynamicesg .
@@
(T3) Rating. A firm’s sustainability performance can be automatically scored to expedite an analyst’s evaluation process. Aggregated rating methods include computing the frequency of sentiment labels for ESG-related headlines fischbach2022automatic , and averaging the E, S, and G topic classification probabilities of documents sokolov2021building . Direct methods explicitly score companies without amalgamating already labelled data. For instance, the prediction of ESG risk ratings from sustainability topics RiskfactorESGpred . Outside of holistic sustainability frameworks (i.e. ESG), direct methods also include estimating carbon emissions from bank transactions.
@@
(T4) Dialogue. Processing sustainability disclosures into interactive dialogue format allows an analyst a user-friendly way for querying information. For instance, a chatbot can be queried to summarise or answer questions about a disclosure Ni2023 .
@@
(T5) Linguistic Patterns. Linguistic analysis deepens an analyst’s understanding of how a disclosure’s style and structure relates with corporate sustainability performance. For example, keyword frequency and word relationships within oil and gas compliance reports can uncover environmental violation patterns Bi2023 , and the syntactic complexity of CSR reports for good CSR performers can also be distinguished Clarkson2020 .
@@ 3 NLP methods for linguistic understanding
 polignano-etal-2022-nlp
 LEE2023119726
 stammbach2023environmental
 BINGLER2022102776 ClimateBERT-NetZero
 Zhou2021
 kang-el-maarouf-2022-finsim4
 finetuneTransformerESG
 tseng2023dynamicesg
 fischbach2022automatic
 sokolov2021building
 RiskfactorESGpred
 gonzalez2022explainable
 mehra2022esgbert Ni2023
 goel2020mining
 Bi2023
 Clarkson2020
 Ni2023
@@ 3.1 Lexical methods
Given their simplicity, lexical methods are appropriate for relatively uncomplicated tasks that do not require significant linguistic understanding. For example, lexical methods can be deployed for task (T5) linguistic patterns, to derive keyword (term) frequency patterns in sustainability disclosures Bi2023 . On the other hand, for more complex tasks such as T1–T4, lexical analysis is utilised as a feature engineering tool rather than a standalone method. For example, TF-IDF vectorisation has been employed to identify companies of interest for ESG classification from news headlines fischbach2022automatic , while TF-IDF-extracted features have been used to detect environmental claims stammbach2023environmental .
@@ 3.2 Semantic methods
Semantic methods interpret textual content to understand its meaning mao2024semantic . This can extend beyond understanding the presence of individual words, to involve contextual understanding and implicit connotations. Unlike lexical methods that simply identify word occurrences, semantic approaches often leverage word embeddings, which are vector representations of words within a multi-dimensional space. We further distinguish these two approaches by analysing the sentence “Renewable energy is sustainable, as it does not utilise finite resources". A lexical approach registers the occurrences of individual words (i.e. ‘renewable’, ‘sustainability’, ‘finite’), without grasping their deeper and interconnected meanings. On the other hand, semantic methods, by determining the proximity of their respective word vectors, interpret ‘renewable’ as synonymous with ‘sustainability’ but antonymous with ‘finite’. This helps semantic algorithms comprehend that ‘renewable’ is conceptually linked to ‘sustainability’, unlike ‘finite’ which implies that resources are exhaustible.
@@
Traditional methods such as GloVe derive text embeddings from word co-occurrences, and can be leveraged to capture semantic relationships pennington2014glove . For instance, similarity methods can be employed on word embeddings to classify sentences according to their relevance to each ESG indicator goel2020mining . However, while useful, these traditional methods may not capture linguistic subtleties comprehensively, paving the way for more advanced techniques.
@@
As a significant leap in semantic understanding, deep learning models such as BERT devlin2018bert powerfully capture the intricacies of text. BERT employs bidirectional attention mechanisms to compute the two-way influence between words. As a result, important contextual and semantic information of a sentence can be captured. BERT develops general language understanding through pre-training on a wide variety of text corpora, and subsequently is fine tuned for specific tasks.
@@
As they attain a high level of semantic understanding, BERT models are suitable for tasks within the corporate sustainability analysis field that require complex language understanding, without being specifically structured for generative tasks. Therefore, BERT is appropriate for complex language understanding tasks like (T1) topic extraction, (T2) sentiment classification and (T3) rating, but not generative tasks such as (T4) dialogue. In line with this, BERT-like models have been used for ESG label classification LEE2023119726 , producing company environmental scores mehra2022esgbert , and to derive topics and their associated keywords for ESG-related risk factors RiskfactorESGpred using BERT-generated text embeddings.
@@
Similar to BERT, generative large language models (Gen-LLMs) such as GPT-4 also involve the comprehension of text by exploiting deep learning based attention architectures achiam2023gpt , and pre-training on large text corpora. However, by learning to predict the next word given the preceding text dong2019unified , their focus extends beyond understanding language, to the generation of coherent text, as well as inference and reasoning capabilities inferencellm . This makes them well-suited for generative answering and summarisation tasks found within (T4) dialogue, and advantageous for reasoning tasks like (T3) rating which involves complex evaluation mao2024gpteval . As such, generative LLMs have been leveraged for transforming TCFD disclosures into dialogue format, as well as evaluating sustainability disclosures for conformity to TCFD reporting guidelines Ni2023 . The latter demonstrates the reasoning abilities of generative LLMs, showing how they can be leveraged to evaluate corporate sustainability performance from disclosures.

### v2

@@ Abstract
journal: Information Fusion
[…]
[1]organization=College of Design and Engineering, National University of Singapore,addressline=9 Engineering Drive 1, postcode=117575, country=Singapore
[…]
[2]organization=Asian Institute of Digital Finance, National University of Singapore,addressline=Innovation 4.0, 3 Research Link, #04-03, postcode=117602, country=Singapore
[…]
[3]organization=College of Computing and Data Science, Nanyang Technological University,addressline=50 Nanyang Ave, postcode=639798, country=Singapore
[…]
[4]organization=Institute of High Performance Computing, Agency for Science, Technology and Research,addressline=Fusionopolis Way, #16-16 Connexis, postcode=138632, country=Singapore
[…]
[5]organization=Sustainable and Green Finance Institute, National University of Singapore,addressline=Innovation 4.0, 3 Research Link, #02-02, postcode=117602, country=Singapore
[…]
[6]organization=NUS Business School, National University of Singapore,addressline=Mochtar Riady Building, 15 Kent Ridge Dr, Mocthar Riady Building, postcode=119245, country=Singapore
[…]
[7]organization=Honorary Research Fellow, Imperial College London,
@@ 1 Introduction
Sustainability, intended as having a non-detrimental impact on the environment, society, and the economy, is becoming increasingly essential for humanity’s future. Various efforts are being undertaken at different levels of the societal and economic hierarchy, from country-level to institutional- and business-level entities, and in some cases, down to individuals. These efforts are focused on making these entities sustainable, amid reputational risks [61], and pressing challenges including global climate change, widespread inequity, governance malpractices, and geopolitical instability [35].
In this work, we focus on sustainability analyses of entities that provide sustainability reports (also referred to as sustainability disclosures); in particular public companies and institutions. These entities are crucial components in the global development of sustainability [14], and we refer to their sustainability as corporate sustainability.
Given these entities’ substantial stake in global sustainability, the analysis of their sustainability, or corporate sustainability analysis, is critical [70]. Yet, this is extremely challenging due to the inherent complexity of corporate sustainability as a concept. To elaborate, the complexities within firms (e.g., geography, size, business activities) and outside the firms (e.g., their global supply chains) as well as the firms’ relationships with non-business stakeholders lead to the evolving frameworks and guidelines put forward to address corporate sustainability issues [11][24]. Further complicating the matter is the lack of globally mandated standards for sustainability reporting that guides companies’ efforts in determining the extent and nature of their sustainability disclosures [5].
For example, a firm may opt to use a convenient corporate sustainability reporting framework that magnifies only certain aspects of a firm’s practices [20], while omitting key sustainability dimensions [23]. Or, within a given regulatory disclosure framework that a firm shall abide to, it may provide sustainability disclosures that are difficult for external parties to digest, and whose integrity and transparency may be questionable [10].
@@
Indeed, today’s corporate sustainability analysis is carried out by human analysts who read sustainability disclosures provided by companies (i.e., the data dimension), and provide corporate sustainability evaluations based on a combination of sustainability frameworks (e.g., Environmental, Social, Governance (ESG) [26], Global Reporting Initiative (GRI) [51], Sustainability Accounting Standards Board (SASB) [30], Greenhouse Gas Protocol and Carbon Disclosure Project (CDP) [65], United Nations’ Sustainable Development Goals (SDG) [67], United Nation Global Compact (UNGC) [62]). These evaluations require an extensive amount of human hours [59], and yet are inevitably influenced by the inherent biases in the frameworks adopted as well as the analysts’ own subjectivity [12][41].
@@
We argue that, to address the subjectivity inherent to data and analysts, it is crucial to adopt artificial intelligence (AI), specifically through linguistic understanding methods enhanced by explainable AI (XAI) capabilities, or what is termed as explainable natural language processing (XNLP) [77]. To qualify, adopting XNLP entails complementing human analysts instead of replacing them. Within corporate sustainability analysis, analysts should remain ‘in the loop’ – while XNLP can expeditiously process numerous sustainability disclosures to unveil consistent and meaningful insights, analysts can further synergise these insights to inform their corporate sustainability assessments.
@@ 2.1 The data dimension
Data is the first building block, or entry point, of corporate sustainability analysis, as depicted in (Fig. 1(A)). While the data required for corporate sustainability assessments can come from various sources, corporate sustainability disclosures are a predominant source [58]. These disclosures consist of reports released by businesses detailing their own sustainability efforts [58]. For example, general sustainability reports typically follow established sustainability frameworks to describe a company’s sustainability efforts and achievements [1], and are typically accompanied by integrated reports that aggregate the general sustainability information with financial implications [32].
@@
Additional data may come from government-mandated sustainability disclosures, where companies are required to release specific aspects of their sustainability performance [80], such as their environmental impact [75], or ESG (Environmental, Social, and Governance) practices [44]. Data originating from NGO sources and media coverage may also be relevant to assess a company’s sustainability efforts [86][4]. Yet these data are influenced by sustainability disclosures [6][33], underlining the importance of focusing our discussion on the sustainability disclosures themselves.
@@
Due to the intricacies and multi-faceted dimensions of corporate sustainability, these data sources inherently contain subjectivity. They involve competing notions of corporate sustainability, expressed through a vast and diverse array of sustainability frameworks [71][57]. Moreover, even within a specific framework, deciding what information to convey necessitates a value judgement (e.g., financial vs. environmental materiality), further exacerbating the partiality involved [7]. We argue that this subjectivity at the data level manifests itself in four data grand challenges that analysts need to face, namely incompleteness, unreliability, ambiguity, and sophistication. These 4 data challenges are reported on the first block of (Fig. 1(A)), referred to as data dimension. We detail them in the following.
@@
Incompleteness. It refers to how data may not provide a comprehensive and complete view of a firm’s sustainability efforts. As certain sustainability disclosures are voluntary, companies can decide on what is material enough to release as sustainability information, thereby raising concerns about openness and the omission of important data [7].
@@
Unreliability. It refers to the trustworthiness and accuracy of sustainability data. Indeed, corporate and government-linked sustainability disclosures may involve greenwashing, where organisations misleadingly claim and exaggerate that their activities or products are more environmentally friendly than they really are [21]. Or they use non-transparent carbon offset products, that often do not deliver the offsets promised [22]. In fact, such developments are particularly frequent for firms that are of larger sizes, to contend with increased dealing with stakeholders [43].
@@
Ambiguity. It refers to data being vague and unclear, making negative content less conspicuous [31]. One example is sharing information without accompanying it with the appropriate context [40]. This may lead analysts to misunderstand the sustainability efforts of companies. Ambiguity is related, to some extent, to greenwashing [91], although it does not explicitly refer to inaccurate or misleading data as unreliability does.
@@
Sophistication. It refers to sustainability disclosures being tediously lengthy, text-heavy, and requiring specialised knowledge to comprehend, making them an extremely sophisticated writing category [76]. As a result, firm sustainability reports can require many hours of human effort to understand, let alone derive useful insights [36].
@@ 2.2 The analyst dimension
Human analysts must synthesise and interpret sustainability data to produce accurate corporate sustainability assessments. To describe this process, relevant data is collected according to the objectives and scope of the corporate sustainability analysis (i.e. evaluating all operations versus one area). Thereafter, the analysis entails leveraging sustainability frameworks or metrics (i.e., carbon footprint) to measure different impact dimensions (environmental, social, economic etc.) [81]. The costs and inaccuracies of this process are exacerbated by human limitations. Specifically, human analysts cannot deal well with the challenges of subjective data, and are not free from subjective interpretations. These human-centric issues are encapsulated by limited human-hours and potential bias.
Limited human-hours. Analysts have a limited amount of time to produce corporate sustainability assessments. However, the process entails extracting meaningful information from corporate sustainability disclosures [36], which requires a significant amount of time i.e., human-hours [59]. In particular, the disclosures are extremely time consuming to read and understand due to the sophistication and ambiguity of the data within them. Moreover, given the potential incompleteness of disclosure information, analysts may analyse additional sources to ascertain a firm’s sustainability, further increasing the human-hours required [68].
Potential Bias. Analysts’ interpretations of sustainability data are subject to bias. This is not least due to the complex and wide ranging factors to be considered, such as firm characteristics, evolving sustainability guidelines, and economic implications, amongst others, often leading to leading to varying interpretations amongst analysts [12][41]. To further compound the potential for bias, data ambiguity and unreliability hinders clear interpretation of sustainability information by obfuscating negative content [31].
@@ 2.3 NLP to the rescue
Subjectivity pervades through the data and analyst dimensions, making corporate sustainability analyses an onerous task. To alleviate this burden, we argue that AI technologies, through linguistic understanding enabled by natural language processing (NLP), can significantly aid analysts. NLP can automatically process hundreds of sustainability disclosures [79], summarising key details and extracting insights for analysts. This reduces the human hours required to analyse disclosures and partially addresses potential bias by providing more consistent and replicable insights [2][17]. We concretise NLP’s usefulness in this domain by detailing five NLP tasks for corporate sustainability analysis (T1-T5) that we deem critical, and that have been partially explored in the literature.
@@
(T1) Topic extraction. Topic extraction allows an analyst to obtain, from sustainability disclosures, content most pertinent to corporate sustainability. Consequently, analysts do not need to sieve through all information, and can focus their analysis on material information [9]. Thematic extraction involves classifying sentences within sustainability frameworks. This includes sorting text into GRI topics [66], ESG pillars [47], or ESG-related concepts [42]. Universal extraction filters textual data into generic categories that are separate from sustainability frameworks. For instance, labelling environmental claims [79], emissions targets [73], or climate relevant information [9]. On the other hand, topic discovery determines topics from sustainability disclosures a posteriori. For example, sustainability topics were uncovered from the sustainability disclosures of shipping companies [92].
@@
(T2) Sentiment Classification. Sentiment insights can also be derived from sustainability disclosures, providing analysts a valuable perspective on corporate sustainability. Polarity classification determines the emotional polarity of text from sustainability disclosures. For instance, within SEC 10k reports, sustainability related sentences can be labelled as positive, negative or neutral [63]. Alternatively, strategic classification uncovers a text’s prospective impact. Examples include the risk and opportunity impact analysis of ESG-related texts [82].
@@
(T3) Rating. A firm’s sustainability performance can be automatically scored to expedite an analyst’s evaluation process. Aggregated rating methods include computing the frequency of sentiment labels for ESG-related headlines [34], and averaging the E, S, and G topic classification probabilities of documents [78]. Direct methods explicitly score companies without amalgamating already labelled data. For instance, the prediction of ESG risk ratings from sustainability topics [46]. Outside of holistic sustainability frameworks (i.e. ESG), direct methods also include estimating carbon emissions from bank transactions.
@@
(T4) Dialogue. Processing sustainability disclosures into interactive dialogue format allows an analyst a user-friendly way for querying information. For instance, a chatbot can be queried to summarise or answer questions about a disclosure [59].
@@
(T5) Linguistic Patterns. Linguistic analysis deepens an analyst’s understanding of how a disclosure’s style and structure relates with corporate sustainability performance. For example, keyword frequency and word relationships within oil and gas compliance reports can uncover environmental violation patterns [8], and the syntactic complexity of CSR reports for good CSR performers can also be distinguished [18].
@@ 3 NLP methods for linguistic understanding
 [66]
 [47]
 [79]
 [9][73]
 [92]
 [42]
 [63]
 [82]
 [34]
 [78]
 [46]
 [37]
 [55][59]
 [36]
 [8]
 [18]
 [59]
@@ 3.1 Lexical methods
Given their simplicity, lexical methods are appropriate for relatively uncomplicated tasks that do not require significant linguistic understanding. For example, lexical methods can be deployed for task (T5) linguistic patterns, to derive keyword (term) frequency patterns in sustainability disclosures [8]. On the other hand, for more complex tasks such as T1–T4, lexical analysis is utilised as a feature engineering tool rather than a standalone method. For example, TF-IDF vectorisation has been employed to identify companies of interest for ESG classification from news headlines [34], while TF-IDF-extracted features have been used to detect environmental claims [79].
@@ 3.2 Semantic methods
Semantic methods interpret textual content to understand its meaning [53]. This can extend beyond understanding the presence of individual words, to involve contextual understanding and implicit connotations. Unlike lexical methods that simply identify word occurrences, semantic approaches often leverage word embeddings, which are vector representations of words within a multi-dimensional space. We further distinguish these two approaches by analysing the sentence “Renewable energy is sustainable, as it does not utilise finite resources”. A lexical approach registers the occurrences of individual words (i.e. ‘renewable’, ‘sustainability’, ‘finite’), without grasping their deeper and interconnected meanings. On the other hand, semantic methods, by determining the proximity of their respective word vectors, interpret ‘renewable’ as synonymous with ‘sustainability’ but antonymous with ‘finite’. This helps semantic algorithms comprehend that ‘renewable’ is conceptually linked to ‘sustainability’, unlike ‘finite’ which implies that resources are exhaustible.
@@
Traditional methods such as GloVe derive text embeddings from word co-occurrences, and can be leveraged to capture semantic relationships [64]. For instance, similarity methods can be employed on word embeddings to classify sentences according to their relevance to each ESG indicator [36]. However, while useful, these traditional methods may not capture linguistic subtleties comprehensively, paving the way for more advanced techniques.
@@
As a significant leap in semantic understanding, deep learning models such as BERT [25] powerfully capture the intricacies of text. BERT employs bidirectional attention mechanisms to compute the two-way influence between words. As a result, important contextual and semantic information of a sentence can be captured. BERT develops general language understanding through pre-training on a wide variety of text corpora, and subsequently is fine tuned for specific tasks.
@@
As they attain a high level of semantic understanding, BERT models are suitable for tasks within the corporate sustainability analysis field that require complex language understanding, without being specifically structured for generative tasks. Therefore, BERT is appropriate for complex language understanding tasks like (T1) topic extraction, (T2) sentiment classification and (T3) rating, but not generative tasks such as (T4) dialogue. In line with this, BERT-like models have been used for ESG label classification [47], producing company environmental scores [55], and to derive topics and their associated keywords for ESG-related risk factors [46] using BERT-generated text embeddings.
@@
Similar to BERT, generative large language models (Gen-LLMs) such as GPT-4 also involve the comprehension of text by exploiting deep learning based attention architectures [3], and pre-training on large text corpora. However, by learning to predict the next word given the preceding text [27], their focus extends beyond understanding language, to the generation of coherent text, as well as inference and reasoning capabilities [72]. This makes them well-suited for generative answering and summarisation tasks found within (T4) dialogue, and advantageous for reasoning tasks like (T3) rating which involves complex evaluation [52]. As such, generative LLMs have been leveraged for transforming TCFD disclosures into dialogue format, as well as evaluating sustainability disclosures for conformity to TCFD reporting guidelines [59]. The latter demonstrates the reasoning abilities of generative LLMs, showing how they can be leveraged to evaluate corporate sustainability performance from disclosures.